OpenStudy (anonymous):

You originally draw a design for an art contest on a 3 in. x 5 in. card. The second phase of the contest requires the drawing to be transferred to an 8.5 in x 11 in. standard sheet of paper and utilize as much of the space on the paper as possible. You determine that the largest size one of the dimensions of your drawing can be is 10.5 in. What is the length of the other dimension if the two drawings are similar? Round to the nearest tenths.

OpenStudy (anonymous):

5/10.5=3/x 5x=31.5 x=6.3 ?idk

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

are you sure it's 3 in. by 5

OpenStudy (anonymous):

yep positive im lookign at the question right now

OpenStudy (anonymous):

preetttyyy suuree its a 3 lol

OpenStudy (anonymous):

all right give me a moment

OpenStudy (anonymous):

and a 5

OpenStudy (anonymous):

you are correct 5/10.5=3/x 5x=31.5 x=6.3

OpenStudy (anonymous):

yes thanks